Express the following statement as a formula with the value of the constant of proportionality determined with the given conditions: w varies directly as x and inversely as the square of y. If x=15 and y=5, then w=36.
Answer by stanbon(75887)

Express the following statement as a formula with the value of the constant of proportionality determined with the given conditions: w varies directly as x and inversely as the square of y. If x=15 and y=5, then w=36.
---------------
w varies directly as x and inversely as the square of y:
w=kx/y
k=wy/x
-------------
If x=15 and y=5, then w=36.
k=36*5/15
k=36/3
k=12
----------
Cheers,
Stan H.